OpenStudy (anonymous):

Simplify: /[2x(2x)^(x-1) + (ln (2x))(2x)^x/] I'm particularly interested in how to work with /[(2x)^(x-1)/]. thanks.

OpenStudy (anonymous):

\[2x(2x)^{x-1}+(Ln(2x))(2x)^{x}\] Is this the question?

OpenStudy (amistre64):

\[a^{(b+c)}=a^b * a^c\]

OpenStudy (anonymous):

They have a natural logarithm there, right?

OpenStudy (amistre64):

yes, but it looks to just care about factoring out the 2x

OpenStudy (anonymous):

Yes, vengeance, sorry about the extra brackets. And yes, amistre64, you gave me the rule I was looking for. Thanks guys. How do I mark these equations so that the system recognizes them as math problems, btw?

OpenStudy (amistre64):

\[(2x)^1*(2x)^{(x-1)}=(2x)^{(x-1+1)}\]

OpenStudy (amistre64):

latex i assume

OpenStudy (anonymous):

@Keketsu=Can you just post the real question please? Lol

OpenStudy (anonymous):

2x(2x)^{x-1} sorry for asking my own question in your question but i just wanted to know with this would you do the factorisation with the outer number first or the indices with the brackets?

OpenStudy (amistre64):

a demonstration to type in latex would be futile since the system would parse it as latex and not a demonstration :)

OpenStudy (anonymous):

@Vengeance921, what you posted is the equation, but it is dealing with the (2x)^(x-1) that I was having trouble with. So that's (2x^x) * (2x^-1) or (2x^x)/2x?

OpenStudy (amistre64):

School, there is no real rule for procedure; only that it makes mathical sense in the end :)

OpenStudy (amistre64):

\[\frac{2x*2x^x}{2x}\]is one way to present it, yes

OpenStudy (anonymous):

Thanks i just saw it in the equation, and i had no idea :P

OpenStudy (anonymous):

Wonderful. Thanks for the help.

OpenStudy (amistre64):

youre welcome